OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x^3y^4)(2x^2y^6)

OpenStudy (anonymous):

please help this is timed!

OpenStudy (anonymous):

6x^5y^10

OpenStudy (saifoo.khan):

she's correct.

OpenStudy (anonymous):

i need a more thorough answer though please help with part 1

OpenStudy (saifoo.khan):

(3x^3y^4)(2x^2y^6) We will multiply the co-efficents with same terms.. while we will solve the variables with variables. (3*2) = 6 For variables, same bases, add exponents, x^3 * x^ 2 = x^(5+2) = x^ 7 y^4 * y^6 = y^(4+6) = y ^ 10